OpenStudy (mathhelp346):

An employee receives a weekly salary of $250 and 8% commission on all sales. Write an equation to find her total weekly earnings if p represents the total amount earned weekly and s represents the total weekly sales. Find her earnings for a week with $1020 sales.

ganeshie8 (ganeshie8):

$250 is fixed salary

ganeshie8 (ganeshie8):

total weekly earnings, p = 250 + 8% of sales

OpenStudy (mathhelp346):

so p=250+.08x?

ganeshie8 (ganeshie8):

p = 250 + 8% of sales sales = s p = 250 + 8% of s p = 250 + 8/100 s p = 250 + .08s

ganeshie8 (ganeshie8):

yes, since they said explicitly sales =s, we call the sales variable s, instead of x :)

OpenStudy (anonymous):

p=331.6

OpenStudy (mathhelp346):

Ok

OpenStudy (anonymous):

Dont forget to give me thumbs up!!